# What Has Been Lost with Synthetic Evaluation?

### v2

@@ 6 Which Benchmark is More Difficult?
Orig. Gen. Orig. Gen. Orig. Gen. Orig. Gen. Orig. Gen.
CondaQA
Original-Data Accuracy 76.7 79.7 69.8 84.1 72.4 83.1 81.0 81.4 70.7 76.3
Full Bundle Consistency 39.5 45.5 37.2 47.7 41.9 52.3 51.2 36.4 39.5 38.6
Paraphrase Consistency 72.7 73.2 70.9 71.4 65.5 78.6 78.2 76.8 67.3 73.2
Scope Consistency 56.3 69.4 47.9 59.2 52.1 67.3 67.3 57.1 47.9 57.1
Affirmative Consistency 57.1 55.8 51.0 61.5 57.1 65.4 51.2 57.7 53.1 50.0
@@
Original-Data Token F1 59.3 81.8 61.7 84.8 74.3 90.9 50.5 74.6 57.9 75.8
Compositional Consistency 28.0 56.0 34.0 60.0 60.0 76.0 28.0 30.0 22.0 44.0
Table 4: Comparison of difficulty of the original CondaQA/DROP samples and their generated version measured by the performance with human-authored questions and edits against generated questions and edits. One author of this paper selected instances with valid generated questions and edits for this comparison. Human annotators provided answers for all of the valid generated questions and edits, and only those with at least 2/3 agreement were used for evaluation. gpt-4-turbo-2024-04-09 gpt-4o-2024-11-20 o3-mini-2025-01-31 meta-llama/Llama-3.3-70B-Instruct Qwen/Qwen2.5-72B-Instruct
@@
Generated questions may be ambiguous. To filter unambiguous for benchmarking, we collect answers from three annotators per question and keep those for which two annotators agree. As NLP expertise is not required here, we recruit 59 annotators through Prolific, all screened using examples with known answers. Screenshots of the answering interfaces are shown in Fig. 12 for CondaQA and Fig. 1317 for DROP. In total, we collect 1242 answers across two datasets at the cost of \$1744.
@@
We use all CondaQA-generated bundles that one author finds valid across the question and all edit types. For benchmarking on original data, we use CondaQA dev questions corresponding to the same passages that have answers in all four passages of their bundle. [^13] If multiple such questions exist, we select one at random. For DROP, we generate questions and compositional edits from passages in the contrast set with human compositional edits, and use question-edit pairs judged valid by one author. For benchmarking on original DROP, we first take human compositional edits from the contrast set corresponding to those passages used for the generated set. We then take the original test set questions from which those edits were created. We report question counts after ambiguity filtering in Table 6 (Appendix). [^14] We manually check if 2 annotators refer to the same span with minor boundary disagreement. If so, the question is considered unambiguous, and one author sets the boundaries.
It has been reported that LLMs exhibit biases when they are used both as a data generator and a task solver Maheshwari et al. (2024), leading to artificially inflated performance when solving evaluation tasks that they themselves have generated. For this reason, we evaluate a range of LLMs from four different model families on both the original and generated datasets: GPT-4-Turbo/GPT-4o, o3-mini, Llama-3.3-70B, and Qwen2.5-72b.
@@
We evaluate CondaQA zero-shot and DROP in a 3-shot setting, following prior work. CondaQA is evaluated using accuracy, as most answers are “yes”, “no”, or “don’t know”. For DROP, we use token-level F1, the harmonic mean of precision and recall over overlapping tokens between the prediction and the gold answer. Both datasets are also evaluated using consistency, which is 1 if all instances in a bundle are answered correctly. For DROP, an instance is considered correct for consistency if its token-level F1 exceeds 0.8. CondaQA bundles consist of a question paired with four passages (“Full Bundle Consistency” in Table 4) or a single passage edited in a specific way. DROP bundles consist of a question and its compositional edit in the context of the same passage.
@@ Results.
They are reported in Table 4, and show that LLMs’ accuracy/token-F1 and consistency are higher in model-generated datasets than in human-authored counterparts, with one exception: Llama-3.3-70B. That is, generated versions are generally easier. We acknowledge that the results for the original DROP sample seem low. This is not due to a bug; we provide a detailed discussion in Appendix D.
@@
Note that this occurs despite model-generated instances often being rated as more valid by humans. This finding suggests that even though generated datasets may closely follow data specifications, and appear to be high-quality to humans, they may feature systematic patterns or lack the creativity of human-authored datasets — thus making them more easily solvable by models. This is evident in part by our finding that human-authored questions in CondaQA tend to be significantly longer than those generated by models (Figure 3).
7 Conclusions
@@
We find that while generated content is often valid (§45), it produces easier benchmarks (§6). Thus, while LLMs are promising for data creation where complexity is less critical, reliable human annotators remain vital for new benchmarks assessing real-world generalization, corner cases, and testing models on nuanced, complex scenarios.
@@ 8 Limitations
First, we focus only on two reasoning-over-text benchmarks. We were constrained by: (i) needing to become deeply familiar with intricate annotation guidelines, (ii) identifying key prompting ideas that ended up being dataset-specific (e.g., reasoning step types for DROP and changing the negation scope through a QA instance generation), (iii) reading passages in these datasets over a thousand times ourselves, (iv) recruiting expert annotators who were either pursuing or had completed a doctoral degree in NLP, (v) designing annotation templates, and (vi) the annotation cost of thousands of dollars.
Second, despite extensive prompting, we may have missed strategies that yield more valid synthetic annotations. Likewise, another LLM might produce more valid generations than gpt-4-turbo-2024-4-09. However, greater validity does not imply greater difficulty, as our comparison of human and generated benchmarks shows. Critically, we first had to discover that valid generations can still be notably easier. Now that is known, improving difficulty would require further rounds of prompt review, costly answer annotation, and manual validity review. A related open question is whether our structured prompts — designed to decompose the task into subtasks with subtask-specific verifiers — improved validity at the cost of difficulty. These prompts were crucial for producing valid generations, yet may have also made the task easier for LLMs. This further highlights the challenge of balancing validity with difficulty.
Finally, in this work, we look at two key metrics for benchmark quality: validity and complexity. However, machine-generated and human-generated benchmarks could differ on several other tertiary aspects which are nevertheless important for building high-quality evaluations, such as diversity, coverage, and representativeness. We leave it to future work to compare these data-creation paradigms on these additional axes.
@@ Acknowledgments
  6. Versions of Table 1 and Table 4, where instead of prompting the model to answer questions in the context of all related passages in a single chat, we use separate chats for each passage-question pair. See: Table 5 and Table 16 (CondaQA only).
  7. Examples of human-authored questions and edits in Tables 7, E, and 9.
@@ Appendix B Contamination Analysis
Our goal is to compare benchmarks with LLM-generated content to their counterparts with human-authored content. But what if generations are actually memorized human instances from the original dataset — a problem know as dataset contamination. This could falsely lead us to conclude that synthetically-generated data is comparable to human-authored data, but these conclusions would then not generalize to new, unseeen datasets.
@@ Appendix E Discussion and Practical Advice
Dataset Context Count
CondaQA All Bundle Passages 44
Original–Paraphrase Passage Pairs 56
Original–Scope Passage Pairs 49
Original–Affirmative Passage Pairs 52
Original Passage 59
DROP Generated Question and Its Edit 50
Generated Questions 66
Compositionally Edited Questions 52
Table 6: The count of generated questions that are unambiguous for all passages in the second column for CondaQA. For DROP, we count generated questions and their edits that are unambiguous; generated questions that are unambiguous (regardless of the edit); and edited questions that are unambiguous (regardless of the question the edit is based on).
@@
Dataset Passage Human-Authored Question Machine-Generated Question
@@
CondaQA The south celestial pole is visible only from the Southern Hemisphere. It lies in the dim constellation Octans, the Octant. Sigma Octantis is identified as the south pole star, more than one degree away from the pole, but with a magnitude of 5.5 it is barely visible on a clear night. Would inexperienced sailors be able to easily use Sigma Octantis to determine their direction of travel within one compass degree of accuracy on a cloudless night? Would Sigma Octantis be considered a reliable star for navigators to use under cloudy skies or in light-polluted areas?
@@
CondaQA It is difficult to distinguish an extinct volcano from a dormant (inactive) one. Dormant volcanoes are those that have not erupted for thousands of years, but are likely to erupt again in the future. Volcanoes are often considered to be extinct if there are no written records of its activity.,. Before its catastrophic eruption of 1991, Pinatubo was an inconspicuous volcano, unknown to most people in the surrounding areas. Two other examples are the … From the way Pinatubo is described in the passage, is it reasonable to assume that it was of fairly modest height, and/or lacking in the usual cone shape that is commonly associated with volcanoes? Would Pinatubo likely have been a main point of interest on a geological tour before its eruption in 1991?
@@
CondaQA Much like elephants, manatees are polyphyodonts, and continuously replace their teeth from the back of the jaw. Adults lack incisors, canines, and premolars, and instead have 8 to 10 cheek teeth in their mouth. Manatees have an infinite supply of teeth moving in from the back and shedding in the front… If the skull of an unknown animal is discovered to have premolars, is it possible that the skull belongs to an adult manatee? If someone discovered a canine tooth, could it belong to an adult manatee?

### v3

@@ 6 Which Benchmark is More Difficult?
GPT-4-Turbo GPT-4o o3-mini Claude-Opus-4 Gemini-2.5-Flash Llama-3.3-70b Qwen2.5-72b
Orig. Gen. Orig. Gen. Orig. Gen. Orig. Gen. Orig. Gen. Orig. Gen. Orig. Gen.
CondaQA
Original-Data Accuracy $73.9_{2.6}$ $\mathbf{79.7_{1.2}}$ $70.5_{0.9}$ $\mathbf{78.3_{1.4}}$ $68.5_{1.9}$ $\mathbf{83.4_{1.4}}$ $67.5_{2.5}$ $\mathbf{80.3_{0.9}}$ $70.2_{0.9}$ $\mathbf{81.7_{1.4}}$ $81.4_{0.0}$ $81.4_{0.0}$ $69.5_{0.0}$ $\mathbf{78.3_{0.8}}$
@@
Paraphrase Consistency $69.5_{1.5}$ $\mathbf{73.9_{1.0}}$ $66.5_{1.6}$ $\mathbf{72.5_{1.0}}$ $65.5_{3.1}$ $\mathbf{78.6_{1.8}}$ $64.7_{1.6}$ $\mathbf{72.1_{3.0}}$ $69.1_{1.8}$ $\mathbf{72.5_{2.7}}$ $\mathbf{78.9_{1.6}}$ $77.1_{1.5}$ $70.2_{1.0}$ $\mathbf{74.6_{1.5}}$
Scope Consistency $53.3_{2.4}$ $\mathbf{67.8_{1.7}}$ $44.2_{0.9}$ $\mathbf{64.5_{3.1}}$ $49.2_{5.6}$ $\mathbf{65.3_{2.0}}$ $49.2_{3.2}$ $\mathbf{62.0_{4.0}}$ $59.6_{2.4}$ $\mathbf{64.5_{1.1}}$ $\mathbf{60.0_{2.7}}$ $58.8_{1.7}$ $49.2_{2.4}$ $\mathbf{59.6_{3.4}}$
Affirmative Consistency $55.9_{4.0}$ $\mathbf{56.2_{2.1}}$ $46.9_{3.2}$ $\mathbf{63.8_{1.6}}$ $51.8_{2.3}$ $\mathbf{61.9_{3.4}}$ $47.8_{2.7}$ $\mathbf{56.2_{3.4}}$ $\mathbf{59.6_{3.4}}$ $53.1_{3.2}$ $\mathbf{68.6_{1.8}}$ $60.0_{0.9}$ $54.7_{1.7}$ $\mathbf{58.1_{2.1}}$
@@
Original-Data Token F1 $61.7_{2.1}$ $\mathbf{83.0_{3.1}}$ $62.8_{2.6}$ $\mathbf{83.9_{1.7}}$ $75.2_{4.1}$ $\mathbf{92.7_{1.3}}$ $73.2_{0.8}$ $\mathbf{90.3_{1.4}}$ $71.9_{1.8}$ $\mathbf{90.9_{0.0}}$ $52.8_{2.0}$ $\mathbf{75.1_{2.8}}$ $57.1_{1.8}$ $\mathbf{77.9_{2.0}}$
Compositional Consistency $30.4_{3.6}$ $\mathbf{58.8_{2.7}}$ $29.2_{3}.0$ $\mathbf{63.6_{0.9}}$ $57.2_{3.0}$ $\mathbf{80.8_{3.0}}$ $55.6_{2.6}$ $\mathbf{82.0_{2.4}}$ $58.0_{2.4}$ $\mathbf{86.4_{1.7}}$ $25.2_{1.1}$ $\mathbf{35.2_{3.3}}$ $23.2_{2.3}$ $\mathbf{48.4_{3.6}}$
Table 4: Comparison of difficulty of the original CondaQA/DROP samples and their generated version measured by the performance with human-authored questions and edits against generated questions and edits. One author of this paper selected instances with valid generated questions and edits for this comparison. Human annotators provided answers for all of the valid generated questions and edits, and only those with at least 2/3 agreement were used for evaluation. Cells highlighted in blue indicate cases where scores on generated data were found to be significantly higher than the scores on the original data, and cells highlighted in yellow indicated cases where scores on original data were found to be significantly higher ($p<.05$). gpt-4-turbo-2024-04-09 gpt-4o-2024-11-20 claude-opus-4-1-20250805 gemini-2.5-flash o3-mini-2025-01-31 meta-llama/Llama-3.3-70B-Instruct Qwen/Qwen2.5-72B-Instruct
@@
Generated questions may be ambiguous. To filter unambiguous questions for benchmarking, we collect answers from three annotators per question and keep those for which two annotators agree. As NLP expertise is not required here, we recruit 59 annotators through Prolific, all screened using examples with known answers. Screenshots of the answering interfaces are shown in Fig. 13 for CondaQA and Fig. 1418 for DROP. In total, we collected 1242 answers across two datasets at a cost of \$1744.
@@
We use all CondaQA-generated bundles that one author finds valid across the question and all edit types. For benchmarking on original data, we use CondaQA dev questions corresponding to the same passages that have answers in all four passages of their bundle. [^14] If multiple such questions exist, we select one at random. For DROP, we generate questions and compositional edits from passages in the contrast set with human compositional edits, and use question-edit pairs judged valid by one author. For benchmarking on original DROP, we first take compositional edits from the human-written dataset for all passages used to create the generated set. We then take the original test set questions from which those human-written compositional edits were created. We report question counts after ambiguity filtering in Table 7 in the Appendix. [^15] We manually check if 2 annotators refer to the same span with minor boundary disagreement. If so, the question is considered unambiguous, and one author sets the boundaries. We provide an illustration of our full synthetic benchmark creation process in Figure 3 in the Appendix.
LLMs can be biased when they are used both as a data generator and a task solver Maheshwari et al. (2024), which causes inflated performance. Thus, we evaluate LLMs from 6 families on the original and generated data: GPT-4-Turbo/GPT-4o [^16] https://openai.com/index/hello-gpt-4o/, o3-mini [^17] https://openai.com/index/openai-o3-mini/, Claude-Opus-4.1181818https://www.anthropic.com/claude/opus, Gemini-2.5-Flash [^19] https://deepmind.google/models/gemini/flash/, Llama-3.3-70B [^20] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, Qwen2.5-72b Team (2024).
@@
We evaluate CondaQA zero-shot and DROP in a 3-shot setting, following prior work. CondaQA is evaluated using accuracy, as most answers are “yes”, “no”, or “don’t know”. For DROP, we use token-level F1, the harmonic mean of precision and recall over overlapping tokens between the prediction and the gold answer. Both datasets are also evaluated using consistency, which is 1 if all instances in a bundle are answered correctly. For DROP, an instance is considered correct for consistency if its token-level F1 exceeds 0.8. CondaQA bundles consist of a question paired with four passages (“Full Bundle Consistency” in Table 4) or a single passage edited in a specific way. DROP bundles consist of a question and its compositional edit in the context of the same passage. We perform 5 runs of the original and generated datasets on each of the models tested. The reported data is the mean value across those 5 runs, and the standard deviation of the mean is shown in subscript. Statistical significance testing is performed to measure the significance of the differences between model performances on the generated and original datasets. We use a mixed effects model to account for variance in the scores across runs, and a one-tailed t-test with $p<.05$ as a threshold for significance.
@@ Results.
They are reported in Table 4, and show that LLMs’ accuracy/token-F1 and consistency tend to be higher in model-generated datasets than in human-authored counterparts. Four out of six model families show higher accuracy/token-F1 and consistency across the board on the generated dataset, with most scores showing a significant increase in model performance. Llama-3.3-70B and Gemini-2.5-Flash on the CondaQA dataset have higher scores on the human-authored datasets, with full bundle consistency and affirmative edit consistency scores on the original dataset being significantly higher than their generated counterparts. Both models have much lower accuracy in answering generated affirmative edits as compared to human-authored edits. Looking into the evaluation logs, we find no discernible pattern as to why generated affirmative edits are harder for these models. However, these are the exceptions to the general trend. Looking at DROP scores, the generated dataset and compositional edits are shown to be significantly easier than the original DROP dataset across all models. That is, generated versions are generally easier. We acknowledge that the results for the original DROP sample seem low. This is not due to a bug; we provide a detailed discussion in Appendix D.
@@
Note that this occurs despite model-generated instances often being rated as more valid by humans. This finding suggests that even though generated datasets may closely follow data specifications, and appear to be high-quality to humans, they may feature systematic patterns or lack the creativity of human-authored datasets — thus making them more easily solvable by models. For example, we observe that the human-authored questions in CondaQA tend to be considerably longer (Figure 4).
7 Conclusion
@@
We find that while generated content is often valid (§45), it tends to produce easier benchmarks (§6). Further, we find that these synthetically-generated datasets do not necessarily preserve model ordering, which could consequently lead practitioners to make different decisions around model development and deployment. Finally, we also find that these quality differences between synthetic and human-authored data, may not be perceptible to researchers inspecting the data. Thus, while LLMs are promising for data creation where complexity is less critical, reliable human annotators remain vital for new benchmarks assessing real-world generalization, corner cases, and testing models on nuanced, complex scenarios.
@@ 8 Limitations
First, we focus only on two reasoning-over-text benchmarks, but seven distinct generation objectives: implications, paraphrasing, changing negation scope, identifying and undoing negation, discrete reasoning, adding intermediate reasoning steps, and altering the temporal order of events. We were constrained by: (i) needing to become deeply familiar with intricate annotation guidelines, (ii) identifying key prompting ideas that ended up being dataset-specific (e.g., reasoning step types for DROP and changing the negation scope through a QA instance generation), (iii) reading passages in these datasets over a thousand times ourselves, (iv) recruiting expert annotators who were either pursuing or had completed a doctoral degree in NLP, (v) designing annotation templates, and (vi) the annotation cost of thousands of dollars.
Second, despite extensive prompting, we may have missed strategies that yield more valid synthetic annotations. Likewise, another LLM might produce more valid generations than gpt-4-turbo-2024-4-09. However, greater validity does not imply greater difficulty, as our comparison of human and generated benchmarks shows. Critically, we first had to discover that valid generations (beyond labels) can still be notably easier. Now that this is known, improving difficulty would require further rounds of prompt review, costly answer annotation, and manual validity review. A related open question is whether our structured prompts — designed to decompose the task into subtasks with subtask-specific verifiers — improved validity at the cost of difficulty. These prompts were crucial for producing valid generations, yet may have also made the task easier for LLMs. This further highlights the challenge of balancing validity with difficulty.
Finally, in this work, we look at two key metrics for benchmark quality: validity and complexity. However, machine-generated and human-generated benchmarks could differ on several other tertiary aspects, which are nevertheless important for building high-quality evaluations, such as diversity, coverage, and representativeness. We leave it to future work to compare these data-creation paradigms on these additional axes.
@@ Acknowledgments
  6. A version of Table 1 and Table 4, where instead of prompting the model to answer questions in the context of all related passages in a single chat, we use separate chats for each passage-question pair. See: Table 5.
  7. Examples of human-authored questions and edits in Tables 6, 8, and 9.
@@
  13. Figure 3: Illustration of our synthetic benchmark creating process. Question and edit generation discussed in §4, label creation discussed in §6.
@@
  16. Examples agreement and disagreement cases between annotators and one author of this paper for generated instances of CondaQA and DROP questions and edits in Tables 16-21.
@@ Appendix B Contamination Analysis
Our goal is to compare benchmarks with LLM-generated content to their counterparts with human-authored content. But what if generations are actually memorized human instances from the original dataset — a problem known as dataset contamination. This could falsely lead us to conclude that synthetically-generated data is comparable to human-authored data, but these conclusions would then not generalize to new, unseeen datasets.
@@ Appendix E Discussion and Practical Advice
Figure 3: Illustration of our synthetic dataset creation process. Purple borders around a block indicates relevance to CondaQA, teal borders indicate relevance to DROP. All other blocks are relevant for both datasets.
@@
Context Count
CondaQA
is-valid(Gen Question, Original Passage) $\wedge$ is-valid(Gen Question, Gen $x$-Edit Passage) 44
$\hookrightarrow\forall x\in\{\text{Paraphrase, Scope, Affirmative}\}$
is-valid(Gen Question, Original Passage) $\wedge$ is-valid(Gen Question, Gen Paraphrase-Edit Passage) 56
is-valid(Gen Question, Original Passage) $\wedge$ is-valid(Gen Question, Gen Scope-Edit Passage) 49
is-valid(Gen Question, Original Passage) $\wedge$ is-valid(Gen Question, Gen Affirmative-Edit Passage) 52
is-valid(Gen Question, Original Passage) 59
DROP
is-valid(Original Passage, Gen Question) $\wedge$ is-valid(Original Passage, Gen Compositional Edit of Gen Question) 50
is-valid(Original Passage, Gen Question) 66
is-valid(Original Passage, Gen Compositional Edit of Generated Question) 52
Table 7: The count of generated instances that are unambiguous. The validity is determined by one of the authors.
@@
Passage Human-Authored Question LLM-Generated Question
@@
The south celestial pole is visible only from the Southern Hemisphere. It lies in the dim constellation Octans, the Octant. Sigma Octantis is identified as the south pole star, more than one degree away from the pole, but with a magnitude of 5.5 it is barely visible on a clear night. Would inexperienced sailors be able to easily use Sigma Octantis to determine their direction of travel within one compass degree of accuracy on a cloudless night? Would Sigma Octantis be considered a reliable star for navigators to use under cloudy skies or in light-polluted areas?
@@
It is difficult to distinguish an extinct volcano from a dormant (inactive) one. Dormant volcanoes are those that have not erupted for thousands of years, but are likely to erupt again in the future. Volcanoes are often considered to be extinct if there are no written records of its activity.,. Before its catastrophic eruption of 1991, Pinatubo was an inconspicuous volcano, unknown to most people in the surrounding areas. Two other examples are […] From the way Pinatubo is described in the passage, is it reasonable to assume that it was of fairly modest height, and/or lacking in the usual cone shape that is commonly associated with volcanoes? Would Pinatubo likely have been a main point of interest on a geological tour before its eruption in 1991?
@@
Much like elephants, manatees are polyphyodonts, and continuously replace their teeth from the back of the jaw. Adults lack incisors, canines, and premolars, and instead have 8 to 10 cheek teeth in their mouth. Manatees have an infinite supply of teeth moving in from the back and shedding in the front […] If the skull of an unknown animal is discovered to have premolars, is it possible that the skull belongs to an adult manatee? If someone discovered a canine tooth, could it belong to an adult manatee?